OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of -2x2 - 5x - 2= 0.

OpenStudy (amistre64):

the discriminate of the quadratic formula is: d = b^2 -4ac if d >0; 2 real roots if d = 0; 1 real root if d<0; no real roots

OpenStudy (amistre64):

ax^2 + bx + c in this case; b=-5 ; a=-2; and c = -2

jhonyy9 (jhonyy9):

D=25-16=9 --- resulted x_1,_2 = (5 +/- 3)/-4 = -2 and -1/2 D>0 --- 2 real solutions